A business to business tech support center is staffed by three representatives.  Calls takes an average of twenty minutes to resolve and calls arrive at a rate of three per hour. The manager is receiving complaints about the waiting time being excessive. If the manager adds two representatives, what is the utilization?
Answer #1

Answer: 0.6

Explanation: number of servers, c= 3+2= 5

mean service rate with 3 servers= 60/20mins= 3 per hour

mu, mean service rate with 1 server= 3/3= 1 per hour

lamda, arrival rate= 3 per hour

Utilization= lamda/(c*mu) = 3/(5*1)= 0.6